OpenStudy (anonymous):

A bat flying in a cave emits a sound and receives its echo 0.1 s later. Show that the distance to the wall of the cave is 17 m. @Abhisar @Sunkist1998

OpenStudy (abhisar):

Ohkay lets begin !

OpenStudy (abhisar):

The sound wave took 0.1 s to reach back to the bat. This means it took only 0.1/2 = 0.05s to reach the wall

OpenStudy (abhisar):

Is it ok ?

OpenStudy (anonymous):

ok

OpenStudy (abhisar):

Now speed of sound in air is 340 m/s

OpenStudy (abhisar):

Distance = Speed * Time

OpenStudy (abhisar):

Distance = 340*0.05=17m

OpenStudy (abhisar):

Got it ?

OpenStudy (anonymous):

yep